Given: x = cost and y = tsint + cost. Find y’.


1
Expert's answer
2022-02-24T11:48:44-0500

It follow that

"y'_x=\frac{y'_t}{x'_t}=\frac{\sin t+t\cos t-\sin t}{-\sin t}=-t\cot t."